# On the quantitative lower bounds for solutions to the Boltzmann equation in nonconvex domains

### v3

@@ 3. Operator estimates and spreading properties
(3.13) $\begin{split}f(t,x,v)&\geq f\left(\frac{t+\tau}{2},x-\frac{t-\tau}{2}v,v\right)e^{-\frac{t-\tau}{2}C_{L}\left<v\right>^{\gamma^{+}}},\\ &\geq\alpha_{m}\left(\tau,\frac{t+\tau}{2},\Delta_{2},A,|v^{\prime}|\right)\mathbf{1}_{B(x^{\prime},\frac{\Delta_{2}}{2^{m}})\times B(v^{\prime},r_{m}(\Delta_{2}))}(x-\frac{t-\tau}{2}v,v)e^{-\frac{t-\tau}{2}C_{L}\left<v\right>^{\gamma^{+}}}\\ &=\alpha_{m}\left(\tau,\frac{t+\tau}{2},\Delta_{2},A,|v^{\prime}|\right)e^{-\frac{t-\tau}{2}C_{L}\left<v\right>^{\gamma^{+}}},\end{split}$
@@
$\geq$ $\alpha_{m}\left(\tau,\frac{t+\tau}{2},\Delta_{2},A,|v^{\prime}|\right)e^{-\frac{t-\tau}{2}C_{L}\left<v\right>^{\gamma^{+}}}\mathbf{1}_{B\left(y,\frac{\Delta_{2}}{2^{m+1}}\right)\times B\left(v^{\prime\prime},\min\{\frac{\Delta_{2}}{2^{m+1}(t-\tau)},1\}\right)}(x,v)$
(3.14) $\geq$ $\alpha_{m}\left(\tau,\frac{\tau^{\prime}+\tau}{2},\Delta_{2},A,|v^{\prime}|\right)e^{-C_{L}\left<\frac{2d_{\Omega}}{\tau^{\prime}-\tau}+1\right>^{\gamma^{+}}}\mathbf{1}_{B\left(y,\frac{\Delta_{2}}{2^{m+1}}\right)\times B\left(v^{\prime\prime},\frac{\Delta_{2}}{2^{m+1}}\right)}(x,v).$
Here, we use Remark 7 in the last inequality above.
The line (3.14) corresponds to (3.12). Thus, we have shown the statement of Proposition 3.3.
@@
Here, the functions $\mathbb{M}(\Delta_{1},\Delta_{2},B,D)$, $\mathbb{B}(m,\tau,\tau^{\prime},\Delta_{2},A)$, $\mathbb{D}(\Delta_{1},\Delta_{2},m)$, $\mathbb{T}(\Delta_{1},\Delta_{2},\tau,m)$, and $\mathbb{V}(\Delta_{1},\Delta_{2},\tau,m,y,x^{\prime})$ are defined as
$\mathbb{M}(\Delta_{1},\Delta_{2},B,D):=\max\left\{1,\left\lceil\log_{\frac{3\sqrt{2}}{4}}\left(\frac{\frac{4d_{\Omega}}{\Delta_{1}}+B+1}{\Delta_{2}}\right)\right\rceil,\left\lceil\log_{2}{\frac{\Delta_{2}+1}{D}}\right\rceil\right\},$
$\mathbb{B}(m,\tau,\Delta_{1},\Delta_{2},A,E):=\alpha_{m}\left(\tau,\tau+\frac{\Delta_{1}}{4},\Delta_{2},A,E\right)e^{-C_{L}\left<\frac{4d_{\Omega}}{\Delta_{1}}+1\right>^{\gamma^{+}}},$
@@
Remark 8.
By Remark 7, we notice that
@@
    $\mathbb{B}(m,\tau,\Delta_{1},\Delta_{2},A,E)$ is strictly decreasing in $E$.
@@
    $\mathbb{M}(\Delta_{1},\Delta_{2},B,D)$ is increasing in $B$ and decreasing in $D$.
@@
Suppose that for fixed $\alpha\in[0,1]$, there exists a continuous mild solution to (1.1)–(1.3). Then, for any $\tau^{\prime\prime}\in(0,\Delta^{0})$, with $\Delta^{0}\in(0,\min\{\frac{d}{10},\frac{1}{2}d(x_{I},\partial\Omega)\}]$ as in Proposition 3.1, there exist positive numbers $\mathfrak{R}(\tau^{\prime\prime})$, $\mathfrak{B}(\tau^{\prime\prime})$, and $\mathfrak{d}(\tau^{\prime\prime})$, depending on $\tau^{\prime\prime}$, $\Omega$, $M$, and $E_{f}$(and $L_{f,p}$ if $\gamma<0$), and a family of velocity vectors
@@
Here, the notation $\mathbb{B}$, $\mathbb{T}$, $\mathbb{D}$, $\mathbb{V}$ are defined in Corollary 3.4. We also define $\mathbb{T}_{0}:=0$ by convention. By Remark 9, we have $\mathbb{T}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0})<\tau^{\prime\prime}$ and that all denominators $\mathbb{T}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0})-\mathbb{T}_{j-1}(\mathfrak{M}_{j}-1,\tau^{\prime\prime},\Delta^{0})$ are positive. We also define
@@
We notice that by Remark 9, we have $\Delta_{1}>0$ and $\Delta_{1}^{\prime}>0$.
@@
Here, we remind that $\mathbb{U}_{1}=\Delta^{0}$. Hence, by noticing that
@@
Hence, we show that (3.21), (3.22), (3.23) holds for $j=k+1$, and we conclude the proof of the statement of Remark 9.
Remark 10.
@@
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the collision kernel $B$ satisfies Assumption B with $\nu<0$, $\alpha\in[0,1]$. Let $f(t,x,v)$ be a continuous mild solution to (1.1)–(1.3). We consider $0<\tau^{\prime\prime}<\Delta^{0}$, where $\Delta^{0}$ is given in Proposition 3.1. Then, there exist $0<\delta_{X}(\tau^{\prime\prime})=\delta_{V}(\tau^{\prime\prime})\leq\frac{d}{56}$, $\delta_{T}(\tau^{\prime\prime})$, $a_{0}(\tau^{\prime\prime})>0$, which depend on $\Omega$, $M$, $E_{f}$, and a modulus of continuity of $f$; and $\mathfrak{R}_{\min}(\tau^{\prime\prime})\geq 2$, which depends on $\tau^{\prime\prime}$, $\Omega$, $M$, and $E_{f}$, such that $\delta_{X}(\tau^{\prime\prime})\geq\delta_{T}(\tau^{\prime\prime})$ and that for all $N\in\mathbb{N}$, there exist $\mathcal{N}(N,\tau^{\prime\prime})$ in $\mathbb{N}$, $\{x_{i}\}_{i=1}^{\mathcal{N}(N,\tau^{\prime\prime})}$ in $\Omega$ and $\{\overline{v}_{i}(\tau^{\prime\prime})\}_{i=1}^{\mathcal{N}(N,\tau^{\prime\prime})}$ in $B(0,\mathfrak{R}_{\min}(\tau^{\prime\prime}))$, such that
@@
Notice that, by the definition and Remark 9, we have
@@
Hence, we have for $\tau^{\prime\prime}\leq t\leq T^{\prime\prime}+\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}+1}\left(\frac{2d_{\Omega}}{T^{\prime\prime}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}+1\right)}$ (notice Remark 7 and the fact that $T^{\prime\prime}+\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}+1}\left(\frac{2d_{\Omega}}{T^{\prime\prime}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}+1\right)}\leq\tau^{\prime\prime}+\mathfrak{d}(\tau^{\prime\prime})$)
@@
(3.50) $\begin{split}&f(t,x,v)\\ \geq&f\left(\frac{t+(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2},x-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v,v\right)e^{-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}C_{L}\left<\mathfrak{R}^{1}_{\min}(\tau^{\prime\prime})+1\right>^{\gamma^{+}}}\\ \geq&\alpha_{\mathfrak{n}^{\prime\prime}}\left(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime})\right)\\ &\mathbf{1}_{B\left(y^{0}_{k(i)},\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}}}\right)\times B(v_{k(i)}(\tau^{\prime\prime}),r_{\mathfrak{n}^{\prime\prime}}(\mathfrak{d}(\tau^{\prime\prime})))}\left(x-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v,v\right)e^{-C_{L}\left<\mathfrak{R}^{1}_{\min}(\tau^{\prime\prime})+1\right>^{\gamma^{+}}}\\ \geq&\alpha_{\mathfrak{n}^{\prime\prime}}\left(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime})\right)e^{-C_{L}\left<\mathfrak{R}^{1}_{\min}(\tau^{\prime\prime})+1\right>^{\gamma^{+}}}.\end{split}$
@@
$f(t,x,v)\geq\alpha_{\mathfrak{n}^{\prime\prime}}\left(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime})\right)e^{-C_{L}\left<\mathfrak{R}^{1}_{\min}(\tau^{\prime\prime})+1\right>^{\gamma^{+}}}$
@@
$\geq\alpha_{\mathfrak{n}^{\prime\prime}}\left(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime})\right)e^{-C_{L}\left<\mathfrak{R}^{1}_{\min}(\tau^{\prime\prime})+1\right>^{\gamma^{+}}}$
@@
$\begin{split}&a_{0}(\tau^{\prime\prime}):=\min\bigg\{\alpha_{\mathfrak{n}^{\prime\prime}}\left(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime})\right)e^{-C_{L}\left<\mathfrak{R}^{1}_{\min}(\tau^{\prime\prime})+1\right>^{\gamma^{+}}},\\ &\hskip 71.13188pt\mathbb{B}(\mathfrak{n}^{\prime\prime},\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime}))\bigg\}.\end{split}$

### v4

@@ 3. Operator estimates and spreading properties
(3.20) $\begin{split}f(t,x,v)&\geq f\left(\frac{t+\tau}{2},x-\frac{t-\tau}{2}v,v\right)e^{-\frac{t-\tau}{2}C_{L}\left\langle v\right\rangle^{\gamma^{+}}},\\ &\geq\alpha_{m}\left(\tau,\frac{t+\tau}{2},\Delta_{2},A,|v^{\prime}|\right)\mathbf{1}_{B\left(x^{\prime},\frac{\Delta_{2}}{2^{m}}\right)\times B(v^{\prime},r_{m}(\Delta_{2}))}(x-\frac{t-\tau}{2}v,v)e^{-\frac{t-\tau}{2}C_{L}\left\langle v\right\rangle^{\gamma^{+}}}\\ &=\alpha_{m}\left(\tau,\frac{t+\tau}{2},\Delta_{2},A,|v^{\prime}|\right)e^{-\frac{t-\tau}{2}C_{L}\left\langle v\right\rangle^{\gamma^{+}}},\end{split}$
@@
$\geq$ $\alpha_{m}\left(\tau,\frac{t+\tau}{2},\Delta_{2},A,|v^{\prime}|\right)e^{-\frac{t-\tau}{2}C_{L}\left\langle v\right\rangle^{\gamma^{+}}}\mathbf{1}_{B\left(y,\frac{\Delta_{2}}{2^{m+1}}\right)\times B\left(v^{\prime\prime},\min\left\{\frac{\Delta_{2}}{2^{m+1}(t-\tau)},1\right\}\right)}(x,v)$
(3.21) $\geq$ $\alpha_{m}\left(\tau,\frac{\tau^{\prime}+\tau}{2},\Delta_{2},A,|v^{\prime}|\right)e^{-C_{L}\left\langle\frac{2d_{\Omega}}{\tau^{\prime}-\tau}+1\right\rangle^{\gamma^{+}}}\mathbf{1}_{B\left(y,\frac{\Delta_{2}}{2^{m+1}}\right)\times B\left(v^{\prime\prime},\frac{\Delta_{2}}{2^{m+1}}\right)}(x,v).$
Here, we use Remark 3.2 in the last inequality above. The line (3.21) corresponds to (3.19). Thus, we have shown the statement of Proposition 3.3.
@@
Here, the functions $\mathbb{M}(\Delta_{1},\Delta_{2},\mathfrak{y}_{1},\mathfrak{y}_{2})$, $\mathbb{B}(m,\tau,\tau^{\prime},\Delta_{2},A,\mathfrak{y}_{3})$, $\mathbb{D}(\Delta_{1},\Delta_{2},m)$, $\mathbb{T}(\Delta_{1},\Delta_{2},\tau,m)$, and $\mathbb{V}(\Delta_{1},\Delta_{2},\tau,m,y,x^{\prime})$ are defined as
$\mathbb{M}(\Delta_{1},\Delta_{2},\mathfrak{y}_{1},\mathfrak{y}_{2}):=\max\left\{1,\left\lceil\log_{\frac{3\sqrt{2}}{4}}\left(\frac{\frac{4d_{\Omega}}{\Delta_{1}}+\mathfrak{y}_{1}+1}{\Delta_{2}}\right)\right\rceil,\left\lceil\log_{2}{\frac{\Delta_{2}+1}{\mathfrak{y}_{2}}}\right\rceil\right\},$
$\mathbb{B}(m,\tau,\Delta_{1},\Delta_{2},A,\mathfrak{y}_{3}):=\alpha_{m}\left(\tau,\tau+\frac{\Delta_{1}}{4},\Delta_{2},A,\mathfrak{y}_{3}\right)e^{-C_{L}\left\langle\frac{4d_{\Omega}}{\Delta_{1}}+1\right\rangle^{\gamma^{+}}},$
@@
Remark 3.3.
By Remark 3.2, we notice that
@@
    $\mathbb{B}(m,\tau,\Delta_{1},\Delta_{2},A,\mathfrak{y}_{3})$ is strictly decreasing in $\mathfrak{y}_{3}$.
@@
    $\mathbb{M}(\Delta_{1},\Delta_{2},\mathfrak{y}_{1},\mathfrak{y}_{2})$ is increasing in $\mathfrak{y}_{1}$ and decreasing in $\mathfrak{y}_{2}$.
@@
Suppose that, for fixed $\alpha\in[0,1]$, there exists a continuous mild solution to (1.1)–(1.3). Then, for any $\tau^{\prime\prime}\in(0,\Delta^{0})$, with $\Delta^{0}\in(0,\min\{\frac{d}{10},\frac{1}{2}d(x_{I},\partial\Omega)\}]$ as in Proposition 3.1, there exist positive numbers $\mathfrak{R}(\tau^{\prime\prime})$, $\mathfrak{B}(\tau^{\prime\prime})$, and $\mathfrak{d}(\tau^{\prime\prime})$, depending on $\tau^{\prime\prime}$, $\Omega$, $M$, and $E_{f}$ (and $L_{f,p}$ if $\gamma<0$), and a family of velocity vectors
@@
Here, the notation $\mathbb{B}$, $\mathbb{T}$, $\mathbb{D}$, $\mathbb{V}$ are defined in Corollary 3.4. We also define $\mathbb{T}_{0}:=0$ by convention. By Remark 3.4, we have $\mathbb{T}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0})<\tau^{\prime\prime}$ and that all denominators $\mathbb{T}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0})-\mathbb{T}_{j-1}(\mathfrak{M}_{j}-1,\tau^{\prime\prime},\Delta^{0})$ are positive. We also define
@@
We notice that by Remark 3.4, we have $\Delta_{1}>0$ and $\Delta_{1}^{\prime}>0$.
@@
Here, we recall that $\mathbb{U}_{1}=\Delta^{0}$. Hence, by noticing that
@@
Hence, we show that (3.28), (3.29), (3.30) holds for $j=k+1$, and we conclude the proof of the statement of Remark 3.4.
Remark 3.5.
@@
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the collision kernel $B$ satisfies Assumption B with $\nu<0$, $\alpha\in[0,1]$. Let $f(t,x,v)$ be a continuous mild solution to (1.1)–(1.3). We consider $0<\tau^{\prime\prime}<\Delta^{0}$, where $\Delta^{0}$ is given in Proposition 3.1. Then, there exist $0<\delta_{X}(\tau^{\prime\prime})=\delta_{V}(\tau^{\prime\prime})\leq\frac{d}{56}$, $\delta_{T}(\tau^{\prime\prime})$, $a_{0}(\tau^{\prime\prime})>0$, which depend on $\Omega$, $M$, $E_{f}$, and a modulus of continuity of $f_{0}$ (and $L_{f,p}$ if $\gamma<0$); and $\mathfrak{R}_{\min}(\tau^{\prime\prime})\geq 2$, which depends on $\tau^{\prime\prime}$, $\Omega$, $M$, and $E_{f}$ (and $L_{f,p}$ if $\gamma<0$), such that $\delta_{X}(\tau^{\prime\prime})\geq\delta_{T}(\tau^{\prime\prime})$ and that for all $N\in\mathbb{N}$, there exist $\mathcal{N}(N,\tau^{\prime\prime})$ in $\mathbb{N}$, $\{x_{i}\}_{i=1}^{\mathcal{N}(N,\tau^{\prime\prime})}$ in $\Omega$ and $\{\overline{v}_{i}(\tau^{\prime\prime})\}_{i=1}^{\mathcal{N}(N,\tau^{\prime\prime})}$ in $B(0,\mathfrak{R}_{\min}(\tau^{\prime\prime}))$, such that
@@
Notice that, by the definition and Remark 3.4, we have
@@
Hence, we have for $\tau^{\prime\prime}\leq t\leq T^{\prime\prime}+\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}+1}\left(\frac{2d_{\Omega}}{T^{\prime\prime}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}+1\right)}$ (see Remark 3.2 and the fact that $T^{\prime\prime}+\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}+1}\left(\frac{2d_{\Omega}}{T^{\prime\prime}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}+1\right)}\leq\tau^{\prime\prime}+\mathfrak{d}(\tau^{\prime\prime})$)
@@
(3.57) $\begin{split}&f(t,x,v)\\ \geq&f\left(\frac{t+(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2},x-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v,v\right)e^{-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}C_{L}\left\langle\mathfrak{R}^{1}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}}\\ \geq&\alpha_{\mathfrak{n}^{\prime\prime}}\left(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime})\right)\\ &\mathbf{1}_{B\left(y^{0}_{k(i)},\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}}}\right)\times B(v_{k(i)}(\tau^{\prime\prime}),r_{\mathfrak{n}^{\prime\prime}}(\mathfrak{d}(\tau^{\prime\prime})))}\left(x-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v,v\right)e^{-C_{L}\left\langle\mathfrak{R}^{1}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}}\\ \geq&\alpha_{\mathfrak{n}^{\prime\prime}}\left(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime})\right)e^{-C_{L}\left\langle\mathfrak{R}^{1}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}}.\end{split}$
@@
$f(t,x,v)\geq\alpha_{\mathfrak{n}^{\prime\prime}}\left(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime})\right)e^{-C_{L}\left\langle\mathfrak{R}^{1}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}}$
@@
$\geq\alpha_{\mathfrak{n}^{\prime\prime}}\left(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime})\right)e^{-C_{L}\left\langle\mathfrak{R}^{1}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}}$
@@
$\begin{split}&a_{0}(\tau^{\prime\prime}):=\min\bigg\{\alpha_{\mathfrak{n}^{\prime\prime}}\left(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime})\right)e^{-C_{L}\left\langle\mathfrak{R}^{1}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}},\\ &\hskip 71.13188pt\mathbb{B}(\mathfrak{n}^{\prime\prime},\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime}))\bigg\}.\end{split}$